OpenStudy (bananas1234):

solve each equation. Write irrational numbers in simplest radical form. Describe the strategy you used to get your solution and tell why you chose that strategy. 1. x^2 + 4 = 0 2. x^2 - 6x + 1 = 0 I think i got number 1. But i need help on question # 2

OpenStudy (anonymous):

How did you work out number 1? Just to be sure... For number two you should use the following algorithm: You have to reform the equation to an equation in like: \[ax^2+bx+c=0\] With a, b and c real numbers. Then you should determine your determinator \[D = b^2-4ac\] And the solutions are: \[x_{1} = \frac{ -b+\sqrt{D} }{ 2a }\] \[x_{2} = \frac{ -b-\sqrt{D} }{ 2a }\]